# Combining Confidence Elicitation and Sample-based Methods for Uncertainty Quantification in Misinformation Mitigation

### 13 Jan 2024 v1

@@ 1 Introduction
It has become crucial to combat the spread of misinformation and detect deceptive content on social media. Misinformation can challenge the fairness of elections (Meel and Vishwakarma, 2020), perpetuate a cascade of rumors resulting in significant financial losses (Marcelo, 2023) and even endanger lives (Loomba et al., 2021). Recent work has demonstrated that Large Language Models (LLMS) can be prime candidates for countering misinformation (Pelrine et al., 2023; Flores and Hao, 2022; Kaliyar et al., 2021; Pelrine et al., 2021). However, their usage in high-value applications is held back by the hallucination problem. The best LLMs have been trained to produce convincing responses, thus they often appear overconfident Ji et al. (2023). Such combination creates instances where the models yield answers that, while sounding reasonable, are significantly inaccurate. Hence, because low uncertainty—or high confidence—does not guarantee accuracy (Huang et al., 2023), it is essential to develop methods to estimate these model’s levels of uncertainty.
@@
We compare the performances of different sample-based consistency methods across various temperature levels and sample sizes. Specifically, we compare several known methods: self-consistency (Wang et al., 2022), an adaptation of selfcheckGPT (Manakul et al., 2023); the normalized standard deviations and the range of predicted class probabilities. We also develop two methods, named SampleAvgDev and Deviation-Sum, and compare their performance with the other sample-based methods. In addition, we explore the distributional and performance shifts of single-step vs. two-step confidence elicitation, showing that the two-step confidence elicitation provides the best calibration. We also carry out comprehensive experiments to evaluate our prompting strategies, including a comparison of the performance of the explain-score prompt (Pelrine et al., 2023) on different truth scales and various versions of GPT. Finally, we integrate all of the above results to the BSDetector framework (Chen and Mueller, 2023), which allows us to evaluate the models’ uncertainty.
@@
  • We compare the calibration capabilities of various sample-based consistency methods in the context of misinformation mitigation, showing how their performances scale with temperature and sample size.
@@
  • We propose the SampleAvgDev sample-based consistency method paired with a two-step confidence elicitation prompt and conclude that this approach is the most efficient calibration technique for our model. attaining an ECE score of 0.076.
@@ 2.1 Misinformation Detection
There are several misinformation detection solutions, which can be categorized into content-based and network-based approaches (Shu et al., 2017). Content-based approaches, the focus of this paper, tackle this issue by analyzing the text, images, or multimedia elements of a message to determine its veracity. While prior solutions’ generalization abilities have been limited (Sharma et al., 2019), GPT-4 has emerged as the top candidate for misinformation detection and classification Pelrine et al. (2023); Quelle and Bovet (2023), demonstrating superior performance on various misinformation datasets. Still, its overall performance and reliability are not robust enough for direct real-world application, as confirmed by Pelrine et al. (2023), which highlight that models often hallucinate and are overconfident in their responses.
@@ 2.2 Uncertainty Quantification
Benefiting from GPT’s impressive verbal capabilities, it is possible to directly elicit these LLM’s uncertainty via verbal cues, such as those demonstrated by Lin et al. (2022) verbalized confidence approach. This technique improves the model’s calibration Tian et al. (2023). Verbalized confidence methods also benefit from prompt engineering principles, where leveraging Chain-of-Thought (CoT) prompting (Wei et al., 2022) enhances the model’s calibration and generates adequate reasoning processes (Xiong et al., 2023).
@@ 3.2 GPT Prompt Engineering
To test GPT’s ability for hard-classification, we provide Politifact’s truth-O-meter scale within the context of the prompt to get a direct 6-point classification truthfulness score. We denote this prompt as ‘Politifact‘. In addition, we tested GPT’s ability for categorical classification on a multiple-choice 3-point scale, i.e., A) True, B) False, C) Somewhat True/ Somewhat False . We denote this prompt as ‘3way-Categorical‘.
@@
To analyze GPT’s distribution and performance shifts in single vs. 2-step confidence elicitation, we prompted the model using different procedures. For the single step confidence elicitation procedure, we request GPT to rate the truthfulness of a LIAR data statement while demanding to rate its uncertainty over that answer, all within a single prompt. We denote this prompt ’single-step-uncertainty’. For the 2-step confidence elicitation procedure, we first obtain a truthfulness score and explanation from the GPT model using the Explain-Score prompt. Then, we prompt the model a second time, now requesting to rate the uncertainty of its previously generated truthfulness score and explanation for the given LIAR data statement. We denote this prompt as ‘2-Step-Uncertainty‘.
@@
Because (CoT) prompting is known to enhances the model’s calibration and generates adequate reasoning processes (Xiong et al., 2023), we devised a prompt inspired from the Explain-Score approach where we specify GPT to generate a truthfulness score paired with a CoT-format explanation. We denote this prompt as ‘CoT-Explain-Score‘.
@@ 3.3 Sample-based consistency methods
In this section, we specify the formulas for the sample-based consistency methods used in our experiments. We assume that the stochastic generated answers $a_{i}$ are from a fixed answer set, $a_{i}\in A$, where $i=1,\ldots,k$ indexes the $k$ candidate outputs sampled. Similarly, we consider an additional non-stochastic reference answer to be $a_{r}$.
@@
Note that we adapted the Self-consistency framework to our context. Specifically, self-consistency’s score is computed as follows:
$a^{*}=\arg\max_{a}\sum_{i=1}^{k}\mathbbm{1}(a_{i}=a)$.
@@
Note that we also adapted the SelfCheckGPT framework to our context. Specifically, selfCheckGPT’s score is computed as follows:
$a^{*}=\frac{1}{k}\ \sum_{i=1}^{k}\mathbbm{1}(a_{i}=a_{r})$.
@@
The sample average deviation (SampleAvgDev) calculates the absolute difference between the sample average and the halfpoint of our classification (50, in our case). Using the halfpoint is motivated by our prompt asking for 100 for something definitively true and 0 for something definitively false (and consequently, maximum uncertainty at the halfpoint). This score is min-max normalized to span the 0-100 range. Specifically, the SampleAvgDev score is computed as follows:
$a^{*}=|\frac{1}{k}\ \sum_{i=1}^{k}(a_{i})-50|$.
@@
The Normalized standard deviation (Norm. std) method involves taking the min-max normalization of the standard deviation of k-stochastic answers.
@@
Deviation-Sum was developed to estimate the model’s uncertainty via the total absolute spread of its answers. Namely, assuming the same mathematical context and $\bar{a_{k}}$ denoting the k-sample average, the Deviation-Sum’s answer is the min-max normalization of the following:
$a^{*}=\sum_{i=1}^{k}|\bar{a_{k}}\ -a_{i}|$.
@@
The predicted class probability range (PredClassMargin) computes the min-max normalization of the margin between the highest and lowest predicted score probabilities, where a wider range implies higher uncertainty, which is particularly relevant to multiclass classification tasks.
@@ 3.4 Evaluation Metrics
This metric is commonly used to evaluate model calibration Tian et al. (2023); Guo et al. (2017). First, we separate the model’s predictions into bins $B_{i}$ with quantile scaling, i.e., the bins are scaled to have the same nmber of examples in each, where $i=1,\ldots,m$ indexes the $m$ bins (we use m = 10). Then, we measure the average accuracy $acc(B_{i})$ and average uncertainty $uncert(B_{i})$ of each bin. Finally, we compute the sum of absolute differences between the average accuracies and uncertainties, weighted by the number of samples $n$ within each bin. A lower ECE implies a better model calibration. Explicitly, the ECE is computed as follows:
$ECE=\sum_{i=i}^{m}\frac{|B_{i}|}{n}\ |acc(B_{i})-uncert(B_{i})|$.
@@
$BrierScore=\frac{1}{N}\ \sum_{i=1}^{N}(uncert(x_{i})-\mathbbm{1}(x_{i}=y_{i}))^{2}$.
@@ 4.2 Effect of sample size
We investigate the effect of varying sampling size on each proposed sample-based consistency method with the premise that the proposed methods might benefit from a larger sample-size. Indeed, previous work has proven that higher sample size leads to better uncertainty estimates in the BSDetector framework (Chen and Mueller, 2023). Table 2 supports this claim, as nearly all proposed methods scale in calibration with sample size. Surprisingly, SampleAvgDev does not require a large sample size to have good calibration. Conversely, Figure 1 displays that selfCheckGPT and PredClassMargin have nearly doubled their improvement in Expected Calibration Error (ECE), implying that the sample size significantly improves the efficacy of these methods.
@@ 4.6 Truth Scales
We also tested with the non-dichotomized 6-way LIAR labels. A challenge here is while we mapped the 0-100 truthfulness scores to the 6-point scale uniformly, Politifact’s Truth-O-meter scale description implies a requirement for a non-uniform mapping. To account for this, we explored different truthfulness scales, and evaluated which scale should be used in the BSDetector Framework. We thus compared the performances of the Explain-Score, Politifact and 3way-Categorical prompts in Table 6 (see Appendix A for a detailed description of each prompt). We see, however, that 6-way performance is quite poor with all approaches, which matches the literature Pelrine et al. (2023) – the 6-way labels may be too subjective, thus, in all the other experiments we focused on the binary ones. In addition, we note that the 3way-Categorical prompt shows poor results.
@@ 5 Conclusion
This study investigated various uncertainty quantification methods to enhance GPT’s ability to provide reliable misinformation mitigation predictions. First, we evaluated different known sample-based consistency methods that capitalized on distinct features of stochastic samples in the context of misinformation mitigation. We demonstrated how each method benefited from high levels of randomness (temperature) and how sample size affected their performance. Then, we confirmed that our two-step confidence elicitation procedure was robust in its calibration and predictive performance. In addition, we explored the robustness of the Explain-Score prompt across different versions of GPT and truthfulness numerical scales. Most importantly, we showed how to successfully implement the BSDetector framework in this domain, which combines sample-based consistency and verbalized methods. In fact, every sample-based consistency method benefited from this hybrid approach, as all proposed methods had better model calibration than their individual counterparts. Overall, our results highlight the superiority of hybrid approaches to uncertainty quantification. Due to the generality of our results, we expect this framework to provide better LLMs model calibration to other misinformation mitigation tasks, moving us closer to a safer application of NLP solutions to combat misinformation.

### 30 Jan 2024 v2

@@ 1 Introduction
It has become crucial to combat the spread of misinformation and detect deceptive content on social media. Misinformation can challenge the fairness of elections (Meel and Vishwakarma, 2020), perpetuate a cascade of rumors resulting in significant financial losses (Marcelo, 2023) and even endanger lives (Loomba et al., 2021). Recent work has demonstrated that Large Language Models (LLMS) can be prime candidates for countering misinformation (Pelrine et al., 2023; Flores and Hao, 2022; Kaliyar et al., 2021; Pelrine et al., 2021). However, their usage in high-value applications is held back by the hallucination problem. The best LLMs have been trained to produce convincing responses, thus they often appear overconfident Ji et al. (2023). Such combination creates instances where the models yield answers that, while sounding reasonable, are significantly inaccurate. Hence, because low uncertainty—or high confidence—does not guarantee accuracy (Huang et al., 2023), it is essential to develop methods to estimate the levels of uncertainty of these models.
@@
We compare the performance of different sample-based consistency methods across various temperature levels and sample sizes. Specifically, we compare several known methods: self-consistency (Wang et al., 2022), an adaptation of selfcheckGPT (Manakul et al., 2023); the normalized standard deviations; and the range of predicted class probabilities. We also develop two methods, named SampleAvgDev and Deviation-Sum, and compare their performance with the other sample-based methods. In addition, we explore the distributional and performance shifts of single-step vs. two-step confidence elicitation, showing that the two-step confidence elicitation provides the best calibration. We also carry out comprehensive experiments to evaluate our prompting strategies, including a comparison of the performance of the explain-score prompt (Pelrine et al., 2023) on different truth scales and various versions of GPT. Finally, we integrate all of the above results to the BSDetector framework (Chen and Mueller, 2023), which allows us to evaluate the models’ uncertainty.
@@
  • We compare the calibration capabilities of various sample-based consistency methods in the context of misinformation mitigation and report how their performances scale with temperature and sample size.
@@
  • We propose the SampleAvgDev sample-based consistency method paired with a two-step confidence elicitation prompt and conclude that this approach is the most efficient calibration technique for our model with an ECE score of 0.076.
@@ 2.1 Misinformation Detection
There are several misinformation detection solutions, which can be categorized into content-based and network-based approaches (Shu et al., 2017). Content-based approaches, the focus of this paper, tackle this issue by analyzing the text, images, or multimedia elements of a message to determine its veracity. While prior solutions’ generalization abilities have been limited (Sharma et al., 2019), GPT-4 has emerged as the top candidate for misinformation detection and classification Pelrine et al. (2023); Quelle and Bovet (2023) by demonstrating superior performance on various misinformation datasets. Still, its overall performance and reliability are not robust enough for direct real-world application, as confirmed by Pelrine et al. (2023), which highlight that models often hallucinate and are overconfident in their responses.
@@ 2.2 Uncertainty Quantification
Benefiting from GPT’s impressive verbal capabilities, it is possible to directly elicit these LLM’s uncertainty via verbal cues, such as those demonstrated by Lin et al. (2022) verbalized confidence approach. This technique improves the model’s calibration Tian et al. (2023). Verbalized confidence methods also benefit from prompt engineering principles, where leveraging Chain-of-Thought (CoT) prompting (Wei et al., 2022) improves the model’s calibration and generates adequate reasoning processes (Xiong et al., 2023).
@@ 3.2 GPT Prompt Engineering
To test GPT’s ability for hard-classification, we provide Politifact’s truth-O-meter scale within the context of the prompt to get a direct 6-point classification truthfulness score. We denote this prompt as ‘Politifact’. In addition, we tested GPT’s ability for categorical classification on a multiple-choice 3-point scale: i.e., A) True, B) False, C) Somewhat True/ Somewhat False. We denote this prompt as ‘3way-Categorical’.
@@
To analyze GPT’s distribution and performance shifts in single vs. 2-step confidence elicitation, we prompted the model using different procedures. For the single step confidence elicitation procedure, we request GPT to rate the truthfulness of a LIAR data statement while demanding to rate its uncertainty over that answer, all within a single prompt. We denote this prompt ‘single-step-uncertainty’. For the 2-step confidence elicitation procedure, we first obtain a truthfulness score and explanation from the GPT model using the Explain-Score prompt. Then, we prompt the model a second time, now requesting to rate the uncertainty of its previously generated truthfulness score and explanation for the given LIAR data statement. We denote this prompt as ‘2-Step-Uncertainty’.
@@
Because (CoT) prompting is known to enhances the model’s calibration and generates adequate reasoning processes (Xiong et al., 2023), we devised a prompt inspired from the Explain-Score approach where we specify GPT to generate a truthfulness score paired with a CoT-format explanation. We denote this prompt as ‘CoT-Explain-Score’.
@@ 3.3 Sample-based consistency methods
In this section, we describe the sample-based consistency methods used in our experiments. For these methods, we generate k-stochastic outputs from the same prompt. We denote the stochastic generated answers $a_{i}$ from a fixed answer set, $a_{i}\in A$, where $i=1,\ldots,k$ indexes the i-th sample. The answer set corresponds to the truthfulness or uncertainty scale used in the prompt. For most of the experiments, A = [0-100]. For selfCheckGPT, we additionally consider an non-stochastic reference answer to be $a_{r}$ generated by setting the parameter T = 0. It is important to note that all sample-based consistency methods were min-max normalized to obtain a common 0-1 uncertainty score.
@@
Note that we attempted to adapt the Self-consistency framework to our context Wang et al. (2022). Specifically, the self-consistency score corresponds to the most frequent score from the k-stochastic answers weighted by $\frac{1}{k}$. It is computed as follows:
$a^{*}=\arg\max_{a}\frac{1}{k}\sum_{i=1}^{k}\mathbbm{1}(a_{i}=a)$
@@
We also attempted to adapt the SelfCheckGPT framework to our context Manakul et al. (2023). Specifically, the selfCheckGPT score is an average of the amount of stochastic answers that match the non-stochastic reference answer. It is computed as follows:
$a^{*}=\frac{1}{k}\ \sum_{i=1}^{k}\mathbbm{1}(a_{i}=a_{r})$
@@
The sample average deviation (SampleAvgDev) calculates the average of the absolute difference between the i-th stochastic answer and the halfpoint of our classification (50, in our case). The rationale behind this method is rooted in our prompt structure: Given that we instruct GPT models to assess the truthfulness of a statement on a 0-100 scale, here 0 represents definitely false and 100 represents definitely true, we can capture the model’s uncertainty by measuring the deviation of its prediction from the halfpoint of our classification (50). Furthermore, averaging these deviations from the halfpoint aims to provide a better representation of the model’s actual uncertainty by the law of large numbers, hence leveraging the principle of consistency for uncertainty quantification. Specifically, the SampleAvgDev score is computed as follows:
$a^{*}=\frac{1}{k}\sum_{i=1}^{k}|a_{i}-50|$
@@
The Normalized standard deviation (Norm. std) method involves taking the standard deviation of k-stochastic answers.
@@
Deviation-Sum was developed to estimate the model’s uncertainty via the total absolute spread of the stochastic answers according to their mean. Namely, letting $\bar{a_{k}}$ denoting the k-sample average, the Deviation-Sum’s answer is computed as follows:
$a^{*}=\sum_{i=1}^{k}|\bar{a_{k}}\ -a_{i}|$
@@
The predicted class probability range (PredClassMargin) computes the margin between the most frequent and the least frequent score. Intuitively, a wider range implies higher uncertainty, and is particularly relevant to multiclass classification tasks.
@@ 3.4 Evaluation Metrics
This metric is commonly used to evaluate model calibration Tian et al. (2023); Guo et al. (2017). First, we separate the model’s predictions into bins $B_{i}$ with quantile scaling, i.e., each bin is scaled to have the same number of examples, where $i=1,\ldots,m$ indexes the $m$ bins (we use m = 10). Then, we measure the average accuracy $acc(B_{i})$ and average uncertainty $uncert(B_{i})$ of each bin. Finally, we compute the sum of absolute differences between the average accuracies and uncertainties, weighted by the number of samples $n$ within each bin. A lower ECE implies a better model calibration. Explicitly, the ECE is computed as follows:
$ECE=\sum_{i=i}^{m}\frac{|B_{i}|}{n}\ |acc(B_{i})-uncert(B_{i})|$
@@
$BrierScore=\frac{1}{N}\ \sum_{i=1}^{N}(uncert(x_{i})-\mathbbm{1}(x_{i}=y_{i}))^{2}$
@@ 4.2 Effect of sample size
We investigate the effect of varying sampling size on each proposed sample-based consistency method with the premise that the proposed methods might benefit from a larger sample-size. Indeed, previous work has proven that higher sample size leads to better uncertainty estimates in the BSDetector framework (Chen and Mueller, 2023). Table 2 supports this claim, as nearly all proposed methods scale in calibration with sample size. Surprisingly, SampleAvgDev does not require a large sample size to have a performing ECE score. Yet, note the decrease in its Brier score suggests it also scales with sample size. Conversely, Figure 1 displays that selfCheckGPT and PredClassMargin have nearly doubled their improvement in Expected Calibration Error (ECE), which suggests that the sample size significantly improves the efficacy of these methods.
@@ 4.6 Truth Scales
We also tested with the non-dichotomized 6-way LIAR labels. A challenge here is while we mapped the 0-100 truthfulness scores to the 6-point scale uniformly, Politifact’s Truth-O-meter scale description implies a requirement for a non-uniform mapping. To account for this, we explored different truthfulness scales, and evaluated which scale should be used in the BSDetector Framework. We thus compared the performances of the Explain-Score, Politifact and 3way-Categorical prompts in Table 6 (see Appendix A for a detailed description of each prompt). We see, however, that 6-way performance is quite poor with all approaches, which matches the literature Pelrine et al. (2023)—the 6-way labels may be too subjective, thus, in all the other experiments we focused on the binary ones. In addition, we note that the 3way-Categorical prompt shows poor results.
@@ 5 Conclusion
This study investigated various uncertainty quantification methods to enhance GPT’s ability to provide reliable misinformation mitigation predictions. First, we evaluated different known sample-based consistency methods that capitalized on distinct features of stochastic samples in the context of misinformation mitigation. We demonstrated how each method benefited from high levels of randomness (temperature) and how sample size affected their performance. Then, we confirmed that our two-step confidence elicitation procedure was robust in its calibration and predictive performance. In addition, we explored the robustness of the Explain-Score prompt across different versions of GPT and truthfulness numerical scales. Most importantly, we showed how to successfully implement the BSDetector framework in this domain, which combines sample-based consistency and verbalized methods. In fact, every sample-based consistency method benefited from this hybrid approach, as all proposed methods had better model calibration than their individual counterparts. Overall, our results highlight the superiority of hybrid approaches to uncertainty quantification. Due to the generality of our results, we expect this framework to provide better LLMs model calibration to other misinformation mitigation tasks and help build safer applications of NLP solutions to combat misinformation.